# LOCAL: Learning with Orientation Matrix to Infer Causal Structure from Time Series Data

### v2

@@ 3.3. Asymptotic Causal Mask Learning
(11) $\sigma_{\tau,\omega}(\boldsymbol{p}_{v}-\boldsymbol{p}_{u})=\frac{exp((\boldsymbol{p}_{v}-\boldsymbol{p}_{u}-\omega)+\dot{g})/\tau}{(exp((\boldsymbol{p}_{v}-\boldsymbol{p}_{u}-\omega)+\dot{g})/\tau)+exp(\ddot{g}/\tau)}$
@@
(12) $\begin{gathered}X_{t}=X_{t}(\boldsymbol{W}\circ\boldsymbol{M}_{\tau,\omega}(\boldsymbol{p})_{u,v})+Y\boldsymbol{A}+Z\\ \end{gathered}$
3.4. Dynamic Graph Parameter Learning
Another problem lies in the existing dynamic causal structure models, which require directly learning $p+1$ $d\times d$-dimensional matrices. There are two main methods to capture dynamic causal effects: 1) Designing a complexity neural network, which specifically customizes a neural network with DAG constraint (Gao et al., 2022; Misiakos et al., 2023). 2) With the help of the Granger causal mechanism, (Sun et al., 2023) introduces the idea of Neural Granger’s component network (Tank et al., 2021) to provide each variable design independent neural network. However, these approaches are quite unintuitive. Overly complex network design or component Neural Granger (Cheng et al., 2024b; Zhou et al., 2024) methods are overparameterized, but most works in practice only exploit the first-layer weights of the network to characterize causal effects, which increases the optimization cost and has to adopt a second-order optimizer L-BFGS-B. Besides, over-parameterization methods are prone to over-fitting when dealing with unsupervised problems such as causal structure discovery, leading to considerable biases in the causal structure.
@@ 3.4. Dynamic Graph Parameter Learning
(13) $\begin{gathered}\boldsymbol{W}=\boldsymbol{E}_{so}(t)\boldsymbol{E}^{T}_{to}(t)\\ \boldsymbol{A}=[\boldsymbol{E}_{so}(t-1)\boldsymbol{E}^{T}_{to}(t-1)|\cdots|\boldsymbol{E}_{so}(t-p)\boldsymbol{E}^{T}_{to}(t-p)]\end{gathered}$
where $\boldsymbol{W}$ is instantaneous dependency, $\boldsymbol{A}$ is lagged dependency. During training, $\boldsymbol{E}_{so}$ and $\boldsymbol{E}_{to}$ will be updated automatically to learn the causal structure among different variables. Compared with the CNN-based work in (Sun et al., 2023), the DGPL module is simpler and the learned embedding matrices have better interpretability (see Section 4.6 for more details). Finally, the DGPL enhanced DBN can be formulated as:
(14) $X_{t}=X_{t}(\boldsymbol{E}_{so}(t)\boldsymbol{E}^{T}_{to}(t)\circ\boldsymbol{M}_{\tau,\omega}(\boldsymbol{p})_{u,v})+Y\boldsymbol{{A}}_{so}\odot\boldsymbol{{A}}_{to}+Z$
@@
3.5. Nonlinear Form and Model Training
3.5.1. Nonlinear Form
@@
(15) $\begin{gathered}X_{t}=\mathcal{W}^{p+1}\ast(\mathcal{W}^{p}\ast\cdots(\mathcal{W}^{1}\ast X_{t-p:t}))=\theta^{T}X_{t-p:t}\\ s.t.\text{ }\mathcal{W}^{c}=\boldsymbol{E}^{c}_{s}(\boldsymbol{E}^{c}_{t})^{T},\forall c\in\{1,\cdots,p+1\}\end{gathered}$
where $\mathcal{W}^{c}$ represents the $c$-th filter, $\boldsymbol{E}^{c}_{so}$ is to learn whether the variable $X_{j}$ is the root cause of other variables in the causal structure, $\boldsymbol{E}^{c}_{to}$ is to learn whether the variable $X_{j}$ is affected by other root causes, while $\boldsymbol{E}_{so}$ and $\boldsymbol{E}_{to}$ are denoted as collections of $\boldsymbol{E}^{c}_{so}$ and $\boldsymbol{E}^{c}_{to}$. $\ast$” is the convolution operation.
Similar to (Sun et al., 2023), The first layer of each CNN is a 1D convolution layer with $p+1$ kernels, stride equal to 1, and no padding, where the last kernels $p+1$ represent instantaneous dependency. Finally, the CNN enhanced LOCAL (LOCAL-CNN) can be formulated as:
(16) $\begin{gathered}\hat{X}_{t}=CNN(X_{t-p:t};\theta^{(q)})\\ s.t.\text{ }\mathcal{W}^{p+1(q)}:=\mathcal{W}^{p+1(q)}\circ\boldsymbol{M}_{\tau,\omega}(\boldsymbol{p}^{(q)})_{u,v})\end{gathered}$
3.5.2. Model Training
@@
(17) $\mathcal{S}(\boldsymbol{W},\boldsymbol{A};X_{t})$ $=\mathcal{S}(\boldsymbol{E}_{so},\boldsymbol{E}_{to},\boldsymbol{p};X_{t})$
@@
(18) $\mathcal{S}(\boldsymbol{W},\boldsymbol{A};X_{t})$ $=\mathcal{S}(\boldsymbol{E}_{so},\boldsymbol{E}_{to},\boldsymbol{p};X_{t})$
@@
4. Experiments
Figure 2. A show case of the result on synthetic data with $d=20$ nodes, $T=1000$ timesteps, $p=1$ lagged order. Our algorithm recovers weights that are close to the ground truth.
In this section, we conduct extensive experiments to answer the following research questions:
  • RQ1: As a framework based on temporal causal discovery, can LOCAL accurately recover causal structures according to the DBN from Eq. (1)?
  • RQ2: Can LOCAL accurately recover causal structures based on complex nonlinear fMRI dataset?
  • RQ3: Can causal structures discovered by LOCAL align with expert knowledge on real-world dataset?
  • RQ4: How do the key designs and parameters in LOCAL affect performance?
  • RQ5: What are the physical meanings of source embedding and target embedding matrices?
4.1. Experimental Setup
@@
The performance of LOCAL is evaluated on three datasets. The first dataset is a synthetic dataset created to test LOCAL’s ability to recover the underlying causal structures. We produce temporal data through a two-step process: 1) sample contemporaneous and time-lagged matrices using the Erdős-Rényi scheme. Populate the contemporaneous matrix $\boldsymbol{W}$ with elements uniformly distributed over the ranges $[-2.0,-0.5]\cup[0.5,2.0]$. For the time-lagged matrices $A_{k}$, sample elements from $[-1.0\alpha,-0.25\alpha]\cup[0.25\alpha,2.0\alpha]$, where $\alpha=1/\eta^{k},\eta\leq 1,k=1,\cdots,p$; 2) generate time series data that adheres to the sampled weighted graph according to the specified equation Eq. (1). The second, NetSim (Smith et al., 2011), is an fMRI dataset designed to facilitate the study of evolving challenges in brain networks. NetSim comprises 28 datasets, from which we select 17 with the same sequence length. Each selected dataset consists of 50 independent time series recordings, capturing the activity of 5-15 nodes over 200 time steps. The third, CausalTime (Cheng et al., 2024b), is a real-world dataset that contains 3 types of benchmark time series from weather, traffic, and healthcare scenarios respectively. The total length of the dataset is 8760 (Air Quality Index subset), 52116 (Traffic subset), and 40000 (Medical subset); the number of nodes is 36 (Air Quality Index subset), 20 (Traffic subset), and 20 (Medical subset).
4.1.2. Hyper-parameter Settings.
@@
In this study, each DGPL module includes a source embedding matrix $\boldsymbol{E}_{so}\in\mathbb{R}^{(P+1)\times d\times k}$ and a target embedding matrix $\boldsymbol{E}_{to}\in\mathbb{R}^{(P+1)\times d\times k}$. For the ACML module, we employ a $\boldsymbol{p}$ vector to represent the priority relationship between nodes and exploit the $\omega$-centered Gumbel sigmoid function to sample the orientation matrix derived. The translation parameter $\omega$ is carefully set to 0.01 to ensure optimal performance. The dimension $k$ of the embedding matrix in DGPL is determined as $\frac{2}{5}\times d$. Under this configuration, the parameter amount of the embedding matrix is marginally smaller than $d\times d$, ensuring efficient representation of the data while mitigating computational complexity. Additionally, the value of the $l_{1}$ penalty term $\lambda_{2}$ is set to 0.01 to regulate the sparsity of the learned causal structures. For synthetic datasets, the batch size is set to 16, while for real-world datasets, a batch size of 64 is utilized. Training and inference were performed on an NVIDIA GeForce RTX 3090 machine, with an average training time of 20 minutes.
4.1.3. Evaluation Metrics
@@
We evaluate the performance of our proposed method for learning causal graphs using three main metrics: 1) True Positive Rate (TPR), which measures the proportion of actual positives that are correctly identified by the model; 2) Structural Hamming Distance (SHD), which counts the number of discrepancies (such as reversed, missing, or redundant edges) between two DAGs; 3) F1 score, which represents the harmonic mean of precision and recall, effectively balancing the two in the evaluation process. Given that the number of potential non-causal relationships vastly outnumbers true causal relationships in real datasets, we utilize the Area Under the Precision-Recall Curve (AUPRC) and the Area Under the ROC Curve (AUROC) to evaluate LOCAL’s effectiveness in identifying genuine and significant causal relationships.
4.1.4. Compared Methods
To answer the first question (RQ1), we compare our LOCAL method with the following popular baselines. Table 1 summarizes the difference between previous methods and LOCAL.
  • DYNOTEARS (Pamfil et al., 2020): DYNOTEARS is a scalable method for learning the structure of DBN from time-series data to efficiently estimate both contemporaneous and time-lagged relationships among variables.
  • NTS-NOTEARS (Sun et al., 2023): NTS-NOTEARS is an extension of the nonlinear CNN version of DYNOTEARS and integrates prior knowledge to improve accuracy.
  • TECDI (Li et al., 2023b): TECDI incorporates intervention data into time series analysis to effectively reveal nonlinear temporal causal relationships, and is the current SOTA model.
Figure 3. TPR, FDR, and F1 scores for different nodes $d=\{5,10,20,50,100\}$. Each row contains results for three different metrics of intra-slice and inter-slice graphs.
Table 1. Difference between existing methods and LOCAL.
Method DYNO NTS TEDCI LOCAL
@@ 4.1.4. Compared Methods
To answer the second question (RQ2), we additionally compare LOCAL with several competitive methods as follows:
  • NGC (Tank et al., 2021): NGC is a neural Granger causal structure learning method that leverages an LSTM or MLP to predict the future and conduct causal discovery based on input weights.
  • PCMCI (Peters et al., 2013): PCMCI is a causal structure learning algorithm based on non-linear independence tests of time series.
To answer the third question (RQ3), we further compare LOCAL with recent or representative causal discovery methods that were directly reproduced by Cheng (Cheng et al., 2024b).
To verify the effectiveness of the quasi-likelihood objective, DGPL module, and ACML module (RQ4), we introduce three variants of LOCAL as follows:
  • LOCAL w/o ACML. It removes the ACML module with a causal constraint.
  • LOCAL w/o DGPL. It removes the DGPL module with a $d\times d$ causal matrix.
  • LOCAL w/o QMLE. It removes the quasi-maximum likelihood estimator (QMLE) with the LSE.
4.2. On Synthetic Datasets (RQ1)
To answer RQ1, we conduct experiments across three simulated datasets, each respectively containing {5, 10, 20, 50, 100} nodes. Each node is accompanied by $p=1$ autoregression terms. Correspondingly, each dataset includes {5, 10, 20, 50, 100} distinct intervention targets, with each target focused on a different node. Figure 2 shows the learned instantaneous matrix $\boldsymbol{W}$ and lagged matrix $\boldsymbol{A}$ on dataset 1 under different algorithms. It can be observed that compared with others, the causal structure learned by LOCAL is almost consistent with the ground truth.
Table 2. Performance on Synthetic Data Under TPR, SHD and F1 score with lower dimension ($d\leq 20$).
@@ 4.2. On Synthetic Datasets (RQ1)
Table 3. Performance on Synthetic Data Under TPR, SHD and F1 score with higher dimension ($d\geq 50$).
@@
Across the synthetic datasets, we conduct three runs of both the baseline methods and LOCAL, averaging the results to obtain final performance indicators. The main results of causal structure learning are shown in Figure 3. To further exploit the effectiveness of LOCAL under the DBN datasets, we also report TPR, SHD, and F1 score results in Table 2 and Table 3. It is evident that when the variable dimension is high (i.e., $d>50$), LOCAL significantly outperforms the baseline model across all three evaluation metrics (TPR, SHD, and F1 score). Notably, LOCAL exhibits a notable improvement of 226.19% in TPR. This superior performance can be attributed to the adaptability of the LOCAL decomposition strategy to high-dimensional data, coupled with the inherently sparse and low-rank structure of the simulation data. Regarding computational efficiency, as depicted in Figure 4, LOCAL demonstrates an average time savings of 97.83% compared to other score-based algorithms. This efficiency gain stems from LOCAL’s avoidance of matrix exponential operations and its utilization of a first-order optimizer. These computational advantages contribute to the practical scalability of LOCAL in analyzing large-scale datasets.
Figure 4. The average running time over 5 datasets measured in minutes with $d=\{5,10,20,50,100\}$, $T=1000$ timesteps, $p=1$ lagged order. The heights of bars are on a logarithmic scale.
4.3. On NetSim Datasets (RQ2)
Table 4. Performance on NetSim Data Under AUROC and AUPRC.

### v3

@@ 2.3 Asymptotic Causal Mask Learning
$\sigma_{\tau,\omega}(\boldsymbol{p}_{v}-\boldsymbol{p}_{u})=\frac{exp((\boldsymbol{p}_{v}-\boldsymbol{p}_{u}-\omega)+\dot{g})/\tau}{(exp((\boldsymbol{p}_{v}-\boldsymbol{p}_{u}-\omega)+\dot{g})/\tau)+exp(\ddot{g}/\tau)}$ (11)
@@
$\begin{gathered}X_{t}=X_{t}(\boldsymbol{W}\circ\boldsymbol{M}_{\tau,\omega}(\boldsymbol{p})_{u,v})+Y\boldsymbol{A}+Z\\ \end{gathered}$ (12)
2.4 Dynamic Graph Parameter Learning
Another problem lies in the existing dynamic causal structure models, which require directly learning $p+1$ $d\times d$-dimensional matrices. There are two main methods to capture dynamic causal effects: 1) Designing a complexity neural network, which specifically customizes a neural network with DAG constraint [18, 30]. 2) With the help of the Granger causal mechanism, [37] introduces the idea of Neural Granger’s component network [38] to provide each variable design independent neural network. However, these approaches are quite counter-intuitive. Overly complex network design or component Neural Granger [12, 44] methods are overparameterized, but most works in practice only exploit the first-layer weights of the network to characterize causal effects, which increases the optimization cost and has to adopt a second-order optimizer L-BFGS-B. Besides, over-parameterization methods are prone to over-fitting when dealing with unsupervised problems such as causal structure discovery, leading to considerable biases in the causal structure.
@@ 2.4 Dynamic Graph Parameter Learning
$\begin{gathered}\boldsymbol{W}=\boldsymbol{E}_{so}(t)\boldsymbol{E}^{T}_{to}(t)\\ \boldsymbol{A}=[\boldsymbol{E}_{so}(t-1)\boldsymbol{E}^{T}_{to}(t-1)|\cdots|\boldsymbol{E}_{so}(t-p)\boldsymbol{E}^{T}_{to}(t-p)]\end{gathered}$ (13)
where $\boldsymbol{W}$ is instantaneous dependency, $\boldsymbol{A}$ is lagged dependency. During training, $\boldsymbol{E}_{so}$ and $\boldsymbol{E}_{to}$ will be updated automatically to learn the causal structure among different variables. Compared with the CNN-based work in [37], the DGPL module is simpler and the learned embedding matrices have better interpretability (see Section 3.6 for more details). Finally, the DGPL enhanced DBN can be formulated as:
$X_{t}=X_{t}(\boldsymbol{E}_{so}(t)\boldsymbol{E}^{T}_{to}(t)\circ\boldsymbol{M}_{\tau,\omega}(\boldsymbol{p})_{u,v})+Y\boldsymbol{{A}}_{so}\odot\boldsymbol{{A}}_{to}+Z$ (14)
@@
2.5 Nonlinear Form and Model Training
2.5.1 Nonlinear Form
@@
$\begin{gathered}X_{t}=\mathcal{W}^{p+1}\ast(\mathcal{W}^{p}\ast\cdots(\mathcal{W}^{1}\ast X_{t-p:t}))=\theta^{T}X_{t-p:t}\\ s.t.\text{ }\mathcal{W}^{c}=\boldsymbol{E}^{c}_{s}(\boldsymbol{E}^{c}_{t})^{T},\forall c\in\{1,\cdots,p+1\}\end{gathered}$ (15)
where $\mathcal{W}^{c}$ represents the $c$-th filter, $\theta^{T}$ represents the weight obtained by combining multiple layers of $\mathcal{W}^{c}$, $\boldsymbol{E}^{c}_{so}$ is to learn whether the variable $X_{j}$ is the root cause of other variables in the causal structure, $\boldsymbol{E}^{c}_{to}$ is to learn whether the variable $X_{j}$ is affected by other root causes, while $\boldsymbol{E}_{so}$ and $\boldsymbol{E}_{to}$ are denoted as collections of $\boldsymbol{E}^{c}_{so}$ and $\boldsymbol{E}^{c}_{to}$. $\ast$” is the convolution operation.
Similar to [37], the first layer of each CNN is a 1D convolution layer with $p+1$ kernels, stride equal to 1, and no padding, where the last kernels $p+1$ represent instantaneous dependency. Finally, the CNN enhanced LOCAL (LOCAL-CNN) can be formulated as:
$\begin{gathered}\hat{X}_{t}=CNN(X_{t-p:t};\theta^{(q)})\\ s.t.\text{ }\mathcal{W}^{p+1(q)}:=\mathcal{W}^{p+1(q)}\circ\boldsymbol{M}_{\tau,\omega}(\boldsymbol{p}^{(q)})_{u,v})\end{gathered}$ (16)
2.5.2 Model Training
@@
$\mathcal{S}(\boldsymbol{W},\boldsymbol{A};X_{t})$ $=\mathcal{S}(\boldsymbol{E}_{so},\boldsymbol{E}_{to},\boldsymbol{p};X_{t})$ (17)
@@
$\mathcal{S}(\boldsymbol{W},\boldsymbol{A};X_{t})$ $=\mathcal{S}(\boldsymbol{E}_{so},\boldsymbol{E}_{to},\boldsymbol{p};X_{t})$ (18)
@@
3 Experiments
3.1 Experimental Setup
@@
The performance of LOCAL was evaluated on three datasets. The first dataset is a synthetic dataset designed to test LOCAL’s ability to recover the underlying causal structures. We generated temporal data through a two-step process: 1) Sampling contemporaneous and time-lagged matrices using the Erdős-Rényi scheme. We populated the contemporaneous matrix $\boldsymbol{W}$ with elements uniformly distributed over the ranges $[-2.0,-0.5]\cup[0.5,2.0]$. For the time-lagged matrices $A_{k}$, elements were sampled from $[-1.0\alpha,-0.25\alpha]\cup[0.25\alpha,2.0\alpha]$, where $\alpha=1/\eta^{k},\eta\leq 1,k=1,\cdots,p$; 2) Generating time series data that adhered to the sampled weighted graph according to Eq. (1). The second dataset, NetSim [36], is an fMRI dataset designed to facilitate the study of evolving challenges in brain networks. NetSim comprises 28 datasets, from which we selected 17 with the same sequence length. Each selected dataset consists of 50 independent time series recordings, capturing the activity of 5-15 nodes over 200 time steps. The third dataset, CausalTime [12], is a real-world benchmark dataset containing three types of time series from weather, traffic, and healthcare scenarios. The total length of the dataset is 8760 (Air Quality Index subset), 52116 (Traffic subset), and 40000 (Medical subset) with node counts of 36, 20, and 20, respectively.
3.1.2 Hyper-parameter Settings.
@@
In this study, each DGPL module included a source embedding matrix $\boldsymbol{E}_{so}\in\mathbb{R}^{(P+1)\times d\times k}$ and a target embedding matrix $\boldsymbol{E}_{to}\in\mathbb{R}^{(P+1)\times d\times k}$. For the ACML module, we employed a $\boldsymbol{p}$ vector to represent the priority relationship between nodes and exploit the $\omega$-centered Gumbel sigmoid function to sample the orientation matrix derived. The translation parameter $\omega$ was carefully set to 0.01 to ensure optimal performance. The dimension $k$ of the embedding matrix in DGPL was determined as $\frac{2}{5}\times d$. Under this configuration, the parameter amount of the embedding matrix was marginally smaller than $d\times d$, ensuring efficient representation of the data while mitigating computational complexity. Additionally, the value of the $l_{1}$ penalty term $\lambda_{2}$ was set to 0.01 to regulate the sparsity of the learned causal structures. For synthetic datasets, the batch size was set to 16, while for real-world datasets, a batch size of 64 was utilized. Training and inference were performed on an NVIDIA GeForce RTX 3090 machine, with an average training time of 20 minutes.
3.1.3 Evaluation Metrics
@@
We evaluated the performance of our proposed method for learning causal graphs using three main metrics: 1) True Positive Rate (TPR), which measured the proportion of actual positives that were correctly identified by the model; 2) Structural Hamming Distance (SHD), which count the number of discrepancies (such as reversed, missing, or redundant edges) between two DAGs; 3) F1 score, which represented the harmonic mean of precision and recall, effectively balancing the two in the evaluation process. Given that the number of potential non-causal relationships vastly outnumbered true causal relationships in real datasets, we utilized the Area Under the Precision-Recall Curve (AUPRC) and the Area Under the ROC Curve (AUROC) to evaluate LOCAL’s effectiveness in identifying genuine and significant causal relationships.
3.1.4 Baselines
Table 1: Comparison of different causal discovery methods. The “Constraint Free” row indicates whether the method imposes constraints on the directed acyclic graph of instantaneous dependencies. The “Augmented Lagrangian” row specifies whether the method employs a Lagrangian augmentation approach. The “Non-linearity” row indicates whether nonlinear dependencies can be captured. The “Mini-batch” row indicates whether the method supports mini-batch training. The “Contemporaneous” row and the “Time-lagged” row refer to the consideration of instantaneous dependencies and time-lagged dependencies.
Method DYNO NTS-NO TECDI LOCAL
@@ 3.1.4 Baselines
To demonstrate the effectiveness of LOCAL in accurately recovering causal structures from Eq. (1) using DBN, we compared it with the following state-of-the-art baselines. Table 1 summarized the difference between previous methods and LOCAL.
  • 1.
    DYNO [34]: DYNO was a scalable method for learning the structure of DBN from time-series data to efficiently estimate both contemporaneous and time-lagged relationships among variables.
  • 2.
    NTS-NO [37]: NTS-NO was an extension of the nonlinear CNN version of DYNO and integrated prior knowledge to improve accuracy.
  • 3.
    TECDI [26]: TECDI incorporated intervention data into time series analysis to effectively reveal nonlinear temporal causal relationships, and was the current SOTA model.
To assess whether LOCAL can accurately recover causal structures from complex nonlinear fMRI datasets, we further compared it with several competitive methods.
  • 1.
    NGC [38]: NGC was a neural Granger causal structure learning method that leveraged an LSTM or MLP to predict the future and conduct causal discovery based on input weights.
  • 2.
    PCMCI [35]: PCMCI was a causal structure learning algorithm based on non-linear independence tests of time series.
To evaluate whether the causal structures identified by LOCAL align with expert knowledge on real-world datasets, we further compared it with recent and representative causal discovery methods, as directly reproduced by Cheng [12].
To verify the effectiveness of the quasi-likelihood objective, DGPL module, and ACML module, we introduced three variants of LOCAL as follows:
  • 1.
    LOCAL w/o ACML. It removed the ACML module with a causal constraint.
  • 2.
    LOCAL w/o DGPL. It removed the DGPL module with a $d\times d$ causal matrix.
  • 3.
    LOCAL w/o QMLE. It removed the quasi-maximum likelihood estimator (QMLE) with the LSE.
3.2 Results on synthetic datasets
Figure 2: Comparison of dependencies learned by various causal discovery models with the ground truth on a synthetic dataset with $d=20$ nodes, $T=1000$ time steps, and a lag order of $p=1$. The results indicate that the dependencies learned by LOCAL closely match the ground truth for both lagged and instantaneous dependencies.
We conducted experiments across three synthetic datasets, each respectively containing {5, 10, 20, 50, 100} nodes. Each node was accompanied by $p=1$ autoregression terms. Correspondingly, each dataset included {5, 10, 20, 50, 100} distinct intervention targets, with each target focused on a different node. Figure 2 showed the learned instantaneous matrix $\boldsymbol{W}$ and lagged matrix $\boldsymbol{A}$ on dataset 1 under different algorithms. It could be observed that compared with others, the causal structure learned by LOCAL was almost consistent with the ground truth.
Figure 3: Comparison of instantaneous dependencies (upper) and lagged dependencies (lower) across various methods on a synthetic dataset with different numbers of nodes, $d=\{5,10,20,50,100\}$. The performance of each method is evaluated using three metrics: True Positive Rate (TPR), Structural Hamming Distance (SHD), and F1 score. Each row present the results for these three metrics, assessing both intra-slice and inter-slice graphs. The results indicate that LOCAL consistently performe well across different metrics and node configurations.
Table 2: Evaluation results of different causal discovery methods on a synthetic dataset with low-dimensional settings ($d\leq 20$). The performance is assessed using True Positive Rate (TPR), Structural Hamming Distance (SHD), and F1 score. Higher TPR and F1 scores indicate better causal discovery performance, while lower SHD values reflect more accurate graph structures.
@@ 3.2 Results on synthetic datasets
Table 3: Evaluation results of different causal discovery methods on a synthetic dataset with high-dimensional settings ($d\geq 50$). The performance is assessed using True Positive Rate (TPR), Structural Hamming Distance (SHD), and F1 score. Higher TPR and F1 scores indicate better causal discovery performance, while lower SHD values reflect more accurate graph structures.
@@
Across the synthetic datasets, we conducted three runs of both the baseline methods and LOCAL, averaging the results to obtain final performance indicators. The main results of causal structure learning were shown in Figure 3. To further exploit the effectiveness of LOCAL under the DBN datasets, we also reported TPR, SHD, and F1 score results in Table 2 and Table 3. It was evident that when the variable dimension was high (e.g., $d>50$), LOCAL significantly outperformed the baseline model across all three evaluation metrics (TPR, SHD, and F1 score). Notably, LOCAL exhibited a notable improvement of 226.19% in TPR. This superior performance could be attributed to the adaptability of the LOCAL decomposition strategy to high-dimensional data, coupled with the inherently sparse and low-rank structure of the simulation data. Regarding computational efficiency, as depicted in Figure 4, LOCAL demonstrated an average time savings of 97.83% compared to other score-based algorithms. This efficiency gained stems from LOCAL’s avoidance of matrix exponential operations and its utilization of a first-order optimizer. These computational advantages contributed to the practical scalability of LOCAL in analyzing large-scale datasets.
Figure 4: Comparison of total runtime and performance across various causal discovery methods. Each method is evaluated on a synthetic dataset with different numbers of nodes, $d=\{5,10,20\}$. The results demonstrate that LOCAL achieve a favorable balance between accuracy and efficiency.
3.3 Results on NetSim Datasets
Table 4: Evaluation results of different causal discovery methods on the NetSim dataset. The performance is assessed using AUROC and AUPRC.